four objects are situated along the y axis as follows: a 2.00 kg object is at +3.00 m,a 3.00-kg object is at 2.50 m,a 2.50 m, a 2.50-kg object is at the origin, and a 400-kg object is at -0.500 m. where is the center of mass of these objects?
1
Expert's answer
2019-12-27T08:57:33-0500

The center of mass of these objects:


"x=\frac{m_1x_1+m_2x_2+m_3x_3+m_4x_4}{m_1+m_2+m_3+m_4}"

"x=\frac{2(3)+3(2.5)+2.5(0)+4(-0.5)}{2+3+2.5+4}=+1.00\ m"